OpenStudy (asiyahh):

The midpoint of line CD is E(-1,0). One endpoint is C(5,2). What are the coordinates of the other end point?

OpenStudy (asiyahh):

@MrNood can you help with this one?

OpenStudy (mrnood):

you know the midpoint - so the distance from one end to the midpoint is the same as the distance from the midpoint to the other end:  So - working always in th same direction work out the change in X from end to midpoint then continue the SAME amount to the other end X and work out the change in Y from end to midpoint then continue the SAME amount to the other end Y

OpenStudy (aum):

One endpoint is C(5,2). Let the other end point be (p, q). The midpoint is given by the formula: ( (p+5)/2, (q+2)/2 ). The midpoint is (-1, 0). Therefore, (p+5)/2 = -1. Solve for p. (q+2)/2 = 0. Solve for q.

OpenStudy (asiyahh):

so would it be (-2.5,0) @aum

OpenStudy (aum):

(p+5)/2 = -1 multiply by 2: p + 5 = -2 add -5: p = -7 (q+2)/2 = 0 multiply by 2: q + 2 = 0 subtract 2: q = -2 The other endpoint is (-7, -2)